OpenStudy (dimples97):

Could someone please help me simplify this expression? -3(b – 7)

MARC (marc_d):

Expand the brackets \[=-3(b-7)\]\[=-3(b)-3(-7)\]

MARC (marc_d):

\[=?\]

MARC (marc_d):

what u get when u multiply -3 with b ?

OpenStudy (dimples97):

3?

MARC (marc_d):

nope

MARC (marc_d):

when u multiply -3 with b, u will get \[=-3\times~b\]\[=-3b\]

OpenStudy (dimples97):

ohhhh okay i see

OpenStudy (dimples97):

my bad

MARC (marc_d):

okay,now what will u get when u multiply -3 with -7? Remember\[(-)negative~value~\times(-)negative~value=(+)positive~value\]\[(+)positive~value~\times~(+)positive~value=(+)positive~value\]\[(-)negative~value~\times~(+)positive~value=(-)negative~value\]\[(+)positive~value~\times(-)negative~value=(-)negative~value\]

MARC (marc_d):

\[=-3\times-7\]\[=?\]

OpenStudy (dimples97):

hmmm 21?

MARC (marc_d):

correct!

OpenStudy (dimples97):

yayyy! lol

MARC (marc_d):

Therefore,the expression will be\[=-3(b-7)\]\[=-3(b)-3(-7)\]\[=-3b+21\]

OpenStudy (dimples97):

Ohhh okay and thats it?? you just solve the last one and thats it?

MARC (marc_d):

yep :)

MARC (marc_d):

@Dimples97

OpenStudy (dimples97):

@MARC_D Thankyouuuu (:

MARC (marc_d):

Yw! @Dimples97 ^_^